# DiffGAD: A Diffusion-based Unsupervised Graph Anomaly Detector

### v1

@@ Abstract
Graph Anomaly Detection (GAD) is crucial for identifying abnormal entities within networks, garnering significant attention across various fields. Traditional unsupervised methods, which decode encoded latent representations of unlabeled data with a reconstruction focus, often fail to capture critical discriminative content, leading to suboptimal anomaly detection. To address these challenges, we present a Diffusion-based Graph Anomaly Detector (DiffGAD). At the heart of DiffGAD is a novel latent space learning paradigm, meticulously designed to enhance the model’s proficiency by guiding it with discriminative content. This innovative approach leverages diffusion sampling to infuse the latent space with discriminative content and introduces a content-preservation mechanism that retains valuable information across different scales, significantly improving the model’s adeptness at identifying anomalies with limited time and space complexity. Our comprehensive evaluation of DiffGAD, conducted on six real-world and large-scale datasets with various metrics, demonstrated its exceptional performance. Our code is available at: https://github.com/fortunato-all/LJH_DiffGAD
@@ 1 Introduction
Graph structure has garnered significant attention from both academia (Wang et al., 2018; Gao et al., 2023c; Roy et al., 2023; 2024; Chen et al., 2023; Gao et al., 2023a; Fang et al., 2024b) and industry (Breuer et al., 2020; Li et al., 2023; Yin et al., 2024; Zhong et al., 2020), with its potential to represent relationships and structures. Among its wide applications, graph anomaly detection (GAD) has evolved as a popular research topic, aiming to detect abnormal targets (nodes, edges, subgraphs, et al.) from the normal.
Existing research can be categorized into two branches: the semi-supervised learning branch (Dou et al., 2020; Liu et al., 2021b; Gao et al., 2023b; 2024b) and the unsupervised learning branch (Hamilton et al., 2017; Chen et al., 2020; Li et al., 2017; Peng et al., 2018). Semi-supervised learning approaches utilize a small subset of labeled data to discern patterns of anomalies, enabling the prediction of labels for the remaining unlabeled dataset. However, human annotation is often time-consuming and labor-intensive, which limits the application of semi-supervised methods.
As alternate, unsupervised reconstruction-based strategies (Ding et al., 2019; Fan et al., 2020; Sakurada & Yairi, 2014) directly capture node characteristics and local structures without the need for annotation. Specifically, these approaches are on the assumption that anomalous entities exhibit more complex distributions and are significantly more difficult to reconstruct (Ding et al., 2019; Fan et al., 2020; Sakurada & Yairi, 2014). Consequently, they utilize encoders to map the graph data into a latent space (Kingma & Welling, 2014; Kipf & Welling, 2017a; Velickovic et al., 2018). Subsequently, the entities are decoded from such latent embeddings, wherein those exhibiting high reconstruction errors are deemed more likely to be anomalies. However, these methods grapple with limitations in their discriminative capability, which results in suboptimal performance. As outlined in (Dou et al., 2020; Liu et al., 2020), abnormal users often deliberately camouflage themselves among normal users. Consequently, they frequently exhibit a significant overlap in common attributes (e.g. age, occupation, length and frequency of reviews, etc.) with normal users. This behavior enables these anomalies to divert the focus of reconstruction-based models towards overlapping common content, reducing the models’ ability to discern truly discriminative features (such as transaction information for fraud detection (Dou et al., 2019)). As a result, anomalous entities are difficult to identify by reconstruction error with such a trivial framework. As illustrated in Figure 1 (a), all data points in the latent space are equally distributed in the learned latent space, without preserving sufficient discriminative content, leading to the overlapping distribution space.
@@
Moreover, given that anomaly detection based on reconstruction error fundamentally aligns more with classification than generation, we introduce slight modifications to the diffusion model to better suit this task. Specifically, during the sampling stage, rather than initiating from a Gaussian distribution, we introduce minor corruptions to the given sample $x$ by adding noises for a small timestep $t<T$, and subsequently denoise it to generate a reconstructed sample $\hat{x}$. Throughout this process, we argue that the general content could be preserved across different scales (reflecting by timestep $t$) to further bolster the model’s discriminative power with confidence. In detail, noise at smaller $t$ values preserves large-scale general content, whereas larger $t$ values maintain more localized general content. As illustrated in Figure 1 (c), by preserving this content, the constructed latent space becomes adept at accurately representing data points. This enhancement effectively widens the distributional discrepancy between anomalous and normal entities, resulting in a robust and discriminative latent space that fosters confident representations.
@@ 2 Background
In this section, we revisit some background. Specifically, we first introduce the preliminaries of the Graph Anomaly Detection (GAD) task, and then describe the diffusion model in latent space.
@@ 3.1 Overview
Figure 2 overviews our proposed DiffGAD, which functions primarily within the latent space. Upon receiving a graph, our methodology initiates by mapping it into the latent space, as detailed in §3.2. Within this transformed domain, we employ a dual Diffusion Model (DM) approach. The first DM is tasked with encapsulating the general content, as elaborated in §3.3, while the second DM targets the extraction of common content, described in §3.4. Herein, the discriminative content is identified by the disparities between these two DM-captured representations. Inspired by (Ho & Salimans, 2022), the concurrent sampling of the two DMs is enabled without necessitating any additional parameters. It is worth noting that our model’s foundation is rooted in a well-established hypothesis, supported by works such as (Ding et al., 2019; Fan et al., 2020; Sakurada & Yairi, 2014), which posits that anomalous entities exhibit more complex distributions and are significantly more difficult to reconstruct. Consequently, nodes with higher reconstruction error are more likely to be anomalies. To facilitate this detection mechanism, our framework requires the original samples with their respective reconstructed counterparts, a process systematically described in §3.3.
@@ 3.3 General Content Preservation
Once the latent space is ready, the reconstruction error by DM acts as a proxy of the anomaly score associated with individual nodes. To achieve this, pairs of original and reconstructed samples are necessary (Gao et al., 2024a). Our general content preservation aims to adapt the DM architecture to the reconstruction error-based anomaly detection task. Unlike traditional DMs, which initiates $\bm{z}_{T}$ as random Gaussian noise to generate $\hat{\bm{z}}_{0}$, our approach perturbs original latent embedding $\bm{z}_{0}$ minimally to get initial points. Concretely, $\bm{z}_{t}$ at a selected small timestep $t<T$ is obtained with Eq. 3, this corrupted embedding is then used to generate reconstructed sample $\hat{\bm{z}}_{0}$ with the aim of preserving the general content. The benefit of this process are threefolds: (1) It facilitates the computation of reconstruction error by aligning original and reconstructed sample; (2) By modulating $t$, the extent of preserved general content can be controlled; a lower $t$ value results in a reconstruction that is closer to the original content, thereby retaining more of the general content. (3) As the DM functions without explicit conditioning, it inherently associates with the general content, which lays the groundwork for the introduction of discriminative content in the following.
@@ 4.2 Ablation Studies
Specifically, we draw the following observations: (1) DiffGAD achieves “an outlier node detection method that works universally well on all datasets”. It obtains top-2 AUC results over all datasets and significantly outperforms current methods with a large margin in average AUC with more than 9% average AUC (compared with the second Average AUC results in ANOMALOUS). (2) DiffGAD demonstrates robustness and stability across different datasets, it attains small std results over different datasets. (3) DiffGAD outperforms other deep-based algorithms by a large margin. Specially, the best improvement is for the Enron dataset, compared with the GAAN method, our DiffGAD attains more than 20.7% AUC gains. These achieved performances are significant and can be attributed to enhancing discriminative ability.
@@ 6 Conclusion
In this work, to tackle the limited ability to capture discriminative content for graph anomaly detection, we make the first effort to transfer the diffusion model from the generation task to a detector and propose a diffusion-based unsupervised graph anomaly detector, namely DiffGAD. Specifically, (1) a discriminative content-guided generation paradigm is proposed to capture the discriminative content and then distill it into the latent space, and (2) a content-preservation strategy is designed to enhance the confidence of the aforementioned guidance process, and (3) extensive experiments on 6 real-world and large-scale datasets with different metrics demonstrate the effectiveness of our method, and (4) comprehensive time and computational analysis demonstrate our efficiency.
Limitation and Future Work. It’s noteworthy that our work is operated in the latent space, but we believe that its expressiveness is limited by the graph encoder, which may restrict the model’s ability to fully capture intricate graph structures and relationships. We will also explore some encoder-free strategies with the help of LLM (Fang et al., 2024a) in the future.
@@ 7 Code of Ethics and Reproducibility Statement
Code of Ethics. This work is primarily foundational in Graph Anomaly Detection, focusing on the development of a more robust, reliable graph anomaly detector. Its primary aim is to contribute to the academic community by enhancing the understanding and implementation of Graph Anomaly Detection techniques. We do not foresee any direct, immediate, or negative societal impacts stemming from the outcomes of our research.
Reproducibility Statement. All results presented in this work are fully reproducible. We provide our code using the anonymous GitHub link. The optimal hyperparameters are detailed in Appendix D.
@@ Appendix A Related Works
Graph Anomaly Detection (GAD) aims to detect abnormal targets from a huge number of normal samples. In the semi-supervised branch, Graph Neural Networks (GNNs) (Gao et al., 2024b; Wang et al., 2019; Liu et al., 2021a; He et al., 2024; Zhang et al., 2024; Chen et al., 2024b; a) have achieved remarkable success in detecting graph anomalies across various domains for its excellent representation capabilities for modeling complex relationships. Specially, CapsGI (Zheng et al., 2024) combines capsules with self-supervised learning (SSL) to overcome the inconsistency of anomaly detection on graphs. PMP (Zhuo et al., 2024) introduces Partitioning Message Passing to adaptively adjust the information aggregated from its heterophilic and homophilic neighbors. RQGNN (Dong et al., 2024) proves that the accumulated spectral energy of the graph signal can be represented by its Rayleigh Quotient, and proposes Rayleigh Quotient Graph Neural Network for graph-level anomaly detection.
To tackle the issue of limited labeled data, recent methods (Fan et al., 2020; Kipf & Welling, 2016; Bandyopadhyay et al., 2020; Yuan et al., 2021; Xu et al., 2022) propose unsupervised approaches over the large-scale unlabeled data, where the anomaly scores is calculated as the reconstruction error. For example, the representative work in (Ding et al., 2019) uses GCN (Kipf & Welling, 2017a) and AE (Kingma & Welling, 2014) to encode the graph into a high-dimensional latent space, decodes the structure and attributes separately using the decoder, and then utilizes the weighted reconstruction error of node feature and structure as the anomaly score.
Diffusion Models (DMs) have garnered widespread attention for their remarkable advances in generating high-quality images and videos. Representative models like DDPM (Ho et al., 2020; Sohl-Dickstein et al., 2015), SMG (Song & Ermon, 2019; 2020), and SDE (Song et al., 2021) have been widely adapted to various domains (Xie et al., 2023; Zhao et al., 2023; Anciukevicius et al., 2023; Bar-Tal et al., 2023; Zhou et al., 2023). For instance, within the scope of GNNs, researchers explore to employ DMs to enhance molecular graph modeling (Huang et al., 2023b; a; Morehead & Cheng, 2023), protein design (Zhang et al., 2023b; Gruver et al., 2023; Wu et al., 2024), drug discovery (Guan et al., 2023; Schneuing et al., 2022), material design (Xie et al., 2022), etc, and achieve significant improvement. There also has been an exploration in GAD, Diga (Li et al., 2023) proposes a semi-supervised DM to detect money laundering, where the DM is guided by labeled anomalies for subgraph recovery. GODM (Liu et al., 2023) introduces a plug-and-play package to adopt a variational encoder and diffusion model to generate effective negative samples to solve class imbalance.
@@ Appendix B Visualization
In this section, we visualize the distribution of the reconstructed representations by using t-SNE (Van der Maaten & Hinton, 2008). For a fair demonstration, we select the smaller-scale Books and the larger-scale Weibo dataset and illustrate the results in Figure 6 and Figure 7, respectively.
@@
Figure 6: Visualization results of the learned representations by (1) reconstruct-based method (e.g. Dominant), and (2) different components of DiffGAD on the Books dataset.
@@ Appendix C More Experimental Details
C.1 Detailed Dataset Descriptions

### v2

@@ Abstract
Graph Anomaly Detection (GAD) is crucial for identifying abnormal entities within networks, garnering significant attention across various fields. Traditional unsupervised methods, which decode encoded latent representations of unlabeled data with a reconstruction focus, often fail to capture critical discriminative content, leading to suboptimal anomaly detection. To address these challenges, we present a Diffusion-based Graph Anomaly Detector (DiffGAD). At the heart of DiffGAD is a novel latent space learning paradigm, meticulously designed to enhance the model’s proficiency by guiding it with discriminative content. This innovative approach leverages diffusion sampling to infuse the latent space with discriminative content and introduces a content-preservation mechanism that retains valuable information across different scales, significantly improving the model’s adeptness at identifying anomalies with limited time and space complexity. Our comprehensive evaluation of DiffGAD, conducted on six real-world and large-scale datasets with various metrics, demonstrated its exceptional performance. Our code is available at: https://github.com/fortunato-all/DiffGAD.
@@ 1 Introduction
Graph structure has garnered significant attention from both academia (Gao et al., 2023c; Roy et al., 2023; Gao et al., 2023a; Fang et al., 2024b) and industry (Breuer et al., 2020; Li et al., 2023; Yin et al., 2024), with its potential to represent relationships and structures. Among its wide applications, graph anomaly detection (GAD) has evolved as a popular research topic, aiming to detect abnormal targets (nodes, edges, subgraphs, et al.) from the normal.
Existing research can be categorized into two branches: the semi-supervised learning branch (Dou et al., 2020; Gao et al., 2023b; 2024b) and the unsupervised learning branch (Hamilton et al., 2017; Chen et al., 2020; Peng et al., 2018). Semi-supervised learning approaches utilize a small subset of labeled data to discern patterns of anomalies, enabling the prediction of labels for the remaining unlabeled dataset. However, human annotation is often time-consuming and labor-intensive, which limits the application of semi-supervised methods.
As alternative, unsupervised strategies (Ding et al., 2019; Fan et al., 2020; Sakurada & Yairi, 2014) directly capture node characteristics and local structures without the need for annotation. Specifically, these approaches are on the assumption that anomalous entities exhibit more complex distributions and are significantly more difficult to reconstruct (Ding et al., 2019; Fan et al., 2020; Sakurada & Yairi, 2014). Consequently, they utilize Autoencoder (AE) to first map the graph data into latent embeddings (Kingma & Welling, 2014; Kipf & Welling, 2017a; Velickovic et al., 2018) and then decoding, wherein those exhibiting high reconstruction errors are deemed more likely to be anomalies. However, these methods grapple with limitations in their discriminative capability, which results in suboptimal performance. As outlined in (Dou et al., 2020; Liu et al., 2020), abnormal users often deliberately camouflage themselves among normal users. Consequently, they frequently exhibit a significant overlap in common attributes (e.g. age, occupation, length and frequency of reviews, etc.) with normal users. This behavior enables these anomalies to divert the focus of reconstruction-based models towards overlapping common content, reducing the models’ ability to discern truly discriminative features (such as transaction information for fraud detection (Dou et al., 2019)). As a result, anomalous entities are difficult to identify by reconstruction error with such a trivial framework. As illustrated in Figure 1 (a), all data points in the latent space are equally distributed in the learned latent space, without preserving sufficient discriminative content, leading to the overlapping distribution space.
@@
Discriminative Distillation to distill discriminative content (Section 3.5).
@@
General Content Preservation to preserve different scale general content (Section 3.3).
Given that anomaly detection based on reconstruction error fundamentally aligns more with classification than generation, we introduce slight modifications to the diffusion model to better suit this task. Specifically, during the sampling stage, rather than initiating from a Gaussian distribution, we introduce minor corruptions to the given sample $x$ by adding noises for a small timestep $t<T$, and subsequently denoise it to generate a reconstructed sample $\hat{x}$. Throughout this process, we argue that the general content could be preserved across different scales (reflecting by timestep $t$) to further bolster the model’s discriminative power with confidence. In detail, noise at smaller $t$ values preserves large-scale general content, whereas larger $t$ values maintain more localized general content. As illustrated in Figure 1 (c), by preserving this content, the constructed latent space becomes adept at accurately representing data points. This enhancement effectively widens the distributional discrepancy between anomalous and normal entities, resulting in a robust and discriminative latent space that fosters confident representations.
@@ 2 Background
In this section, we revisit some background. Specifically, we first introduce the preliminaries of the Graph Anomaly Detection (GAD) task and then describe the diffusion model in latent space.
@@ 3.1 Overview
Figure 2 overviews our proposed DiffGAD, which functions primarily within the latent space. Upon receiving a graph, our methodology initiates by mapping it into the latent space, as detailed in §3.2. Within this transformed domain, we employ a dual Diffusion Model (DM) approach. The first DM is tasked with encapsulating the general content, as elaborated in §3.3, while the second DM targets the extraction of common content, described in §3.4. Herein, the discriminative content is identified by the disparities between these two DM-captured representations. Inspired by (Ho & Salimans, 2022), the concurrent sampling of the two DMs can be enabled by controlling a simple hyper-parameter. It is worth noting that our model’s foundation is rooted in a well-established hypothesis, supported by works such as (Ding et al., 2019; Fan et al., 2020; Sakurada & Yairi, 2014), which posits that anomalous entities exhibit more complex distributions and are significantly more difficult to reconstruct. Consequently, nodes with higher reconstruction error are more likely to be anomalies. To facilitate this detection mechanism, our framework requires the original samples with their respective reconstructed counterparts, a process systematically described in §3.3.
@@ 3.3 General Content Preservation
Once the latent space is ready, the reconstruction error by DM acts as a proxy of the anomaly score associated with individual nodes. To achieve this, pairs of original and reconstructed samples are necessary (Gao et al., 2024a). Our general content preservation aims to adapt the DM architecture to the reconstruction error-based anomaly detection task. Unlike traditional DMs, which samples random Gaussian noise to generate $\hat{\bm{z}}_{0}$, our approach perturbs original latent embedding $\bm{z}_{0}$ by mixing random noises from different scales as the initial points for sampling. Concretely, following the work in Ho et al. (2020), we categorize the noises into 500 different scales, and we add noises at each scale by the following equation:
$z_{t}=\frac{t}{T}z_{0}+\frac{T-t}{T}\epsilon$ (8)
where $\epsilon$ is a random noise, and $\epsilon\in\mathcal{N}(0,1)$, this corrupted embedding $z_{t}$ is then used to generate reconstructed sample $\hat{\bm{z}}_{0}$ with the aim of preserving the general content. The benefit of this process are threefolds: (1) It facilitates the computation of reconstruction error by aligning original and reconstructed sample; (2) By modulating $t$, the extent of preserved general content can be controlled; a lower $t$ value results in a reconstruction that is closer to the original content, thereby retaining more of the general content. (3) As the DM functions without explicit conditioning, it inherently associates with the general content, which lays the groundwork for the introduction of discriminative content.
@@ 4.2 Ablation Studies
Specifically, we draw the following observations: (1) DiffGAD achieves “an outlier node detection method that works universally well on all datasets”. It obtains top-2 AUC results over all datasets and significantly outperforms current methods with a large margin in average AUC with more than 9% average AUC (compared with the second Average AUC results in ANOMALOUS). (2) DiffGAD demonstrates robustness and stability across different datasets, it attains small std results over different datasets. (3) DiffGAD outperforms other deep-based algorithms by a large margin. Specially, the best improvement is for the Enron dataset, compared with the GAAN method, our DiffGAD attains more than 20.7% AUC gains. These achieved performances are significant, and this can be attributed to the enhancement of discriminative ability. More detailed analysis is in Appendix C.1.
@@ 6 Conclusion
In this work, we make the first effort to transfer the diffusion model from generation task to a detector and propose a diffusion-based unsupervised graph anomaly detector, namely DiffGAD. Specifically, (1) a discriminative content-guided generation paradigm is proposed to capture the discriminative content and then distill it into the latent space, and (2) a content-preservation strategy is designed to enhance the confidence of the aforementioned guidance process, and (3) extensive experiments on 6 real-world and large-scale datasets with different metrics demonstrate the effectiveness of our method, and (4) comprehensive time and computational analysis demonstrate our efficiency.
Limitation and Future Work. The expressiveness of latent embeddings might be limited by graph encoder, we will explore some encoder-free strategies in the future.
@@ 7 Code of Ethics and Reproducibility Statement
Code of Ethics. We do not foresee any direct, immediate, or negative societal impacts stemming from the outcomes of our research. Reproducibility Statement. All results presented in this work are fully reproducible. We provide our code using the GitHub link. The optimal hyperparameters are detailed in Appendix D.
8 Acknowledgements
This research is supported by the National Natural Science Foundation of China (92270114). This research was also supported by the advanced computing resources provided by the Supercomputing Center of the USTC.
@@ Appendix A Related Works
Graph Anomaly Detection (GAD) aims to detect abnormal targets from a huge number of normal samples. In the semi-supervised branch, Graph Neural Networks (GNNs) (Wang et al., 2019; He et al., 2024; Zhang et al., 2024) have achieved remarkable success in detecting graph anomalies across various domains for its excellent representation capabilities for modeling complex relationships. Specially, CapsGI (Zheng et al., 2024) combines capsules with self-supervised learning (SSL) to overcome the inconsistency of anomaly detection on graphs. PMP (Zhuo et al., 2024) introduces Partitioning Message Passing to adaptively adjust the information aggregated from its heterophilic and homophilic neighbors. RQGNN (Dong et al., 2024) proves that the accumulated spectral energy of the graph signal can be represented by its Rayleigh Quotient, and proposes Rayleigh Quotient GNN for graph-level anomaly detection.
To tackle the issue of limited labeled data, recent methods (Fan et al., 2020; Yuan et al., 2021; Xu et al., 2022) propose unsupervised approaches over the large-scale unlabeled data, where the anomaly score is calculated as the reconstruction error. For example, the representative work in (Ding et al., 2019) uses GCN (Kipf & Welling, 2017a) and AE (Kingma & Welling, 2014) to encode the graph into a high-dimensional latent space, decodes the structure and attributes separately using the decoder, and then utilizes the weighted reconstruction error of node feature and structure as the anomaly score.
Diffusion Models (DMs) have garnered widespread attention for their remarkable advances in generating high-quality images and videos. Fueled by the explosion of deep learning (Lu et al., 2025; Fang et al., 2024a; Lu et al., 2023; 2024), representative models like DDPM (Ho et al., 2020; Sohl-Dickstein et al., 2015), SMG (Song & Ermon, 2019; 2020), and SDE (Song et al., 2021) have been widely adapted to various domains (Xie et al., 2023; Mao et al., 2025). For instance, within the scope of GNNs, researchers explore to employ DMs to enhance molecular graph modeling (Huang et al., 2023b; a), protein design (Zhang et al., 2023b; Gruver et al., 2023), drug discovery (Guan et al., 2023; Schneuing et al., 2022), material design (Xie et al., 2022), etc, and achieve significant improvement. There also has been an exploration in GAD, Diga (Li et al., 2023) proposes a semi-supervised DM to detect money laundering, where the DM is guided by labeled anomalies for subgraph recovery. GODM (Liu et al., 2023) introduces a plug-and-play package to adopt a variational encoder and diffusion model to generate effective negative samples to solve the class imbalance.
@@ Appendix B Visualization
In this section, we visualize the distribution of the reconstructed representations by using t-SNE (Van der Maaten & Hinton, 2008). For a fair demonstration, we select the smaller-scale Books and the larger-scale Weibo dataset and illustrate the results in Figure 6 and Figure 7, respectively. Moreover, for clearer demonstration, we refine the visualization of Books in Figure 8.
@@
Figure 6: Visualization results of the learned representations by (1) reconstruct-based method (e.g. Dominant), and (2) different components of DiffGAD on the Books dataset. Variances of points in Graph Auto Encoder: [6.066866e-22, 3.768135e-35], very close to 0.
@@
Figure 8: Visualization results of the learned representations by different components of DiffGAD on the Books dataset, for clear illustration, we construct the boundary by connecting neighbors and highlight the abnormal samples. Where (a), (b), (c) are with small t (t=100), and (d), (e), (f) are with large t (t=400), please zoom in for clearer observation.
@@ Appendix C More Experimental Details
C.1 Detailed Experimental Analysis
In this subsection, we analyze the experimental results of our DiffGAD over the datasets where we achieve sub-optimal performance.
Weibo. The success of most methods on Weibo is because the outliers in Weibo exhibit the properties of both structural and contextual outliers. Specifically, in Weibo, the average clustering coefficient of the outliers is higher than that of inliers (0.400 vs. 0.301), meaning that these outliers correspond to structural outliers. Meanwhile, the average neighbor feature similarity of the outliers is far lower than that of inliers (0.004 vs. 0.993), so the outliers also correspond to contextual outliers.
The two classical algorithms Radar and ANOMALOUS employ graph-based features to extract the anomalous signals from graphs to more flexibly encode graph information to spot outlier nodes, which helps them perform best on Weibo. As they constraint on graph/node types or prior knowledge, they can not generalize well on other datasets.
Reddit & Dgraph. In contrast, the outliers in the Reddit and DGraph datasets have similar average neighbor feature similarities and clustering coefficients for outliers and inliers. Therefore, their abnormalities rely more on outlier annotations with domain knowledge.
Non-graph algorithm LOF (Local Outlier Factor) identifies local outliers based on density while IF (Isolation Forest) builds an ensemble of base trees to isolate the data points and defines the decision boundary as the closeness of an individual instance to the root of the tree. Both of them solely use node attributes thus avoiding the influence of structures.
Disney. DiffGAD and most of the deep algorithms do not work particularly well on Disney compared to classical baselines. The reason is that Disney has small graphs in terms of ‘Nodes’, ‘Edges’, and ‘Features’ (See Table 5 ). The small amount of data could make it difficult for the deep learning methods to encode the inlier distribution well and could also possibly lead to overfitting issues.
C.2 Detailed Dataset Descriptions